# Equiripple MIMO Beampattern Synthesis using Chebyshev Approximation

### v1

@@ I Introduction
The majority of the efforts in the literature focus on designing beampatterns with a nominally flat passband region, a narrow transition band region, and a stopband region where almost no power is transmitted. As a consequence of the Gibbs phenonmenon, the beampattern’s passband stopband regions possess ripples. The magnitude of these ripples appears to be controlled by the width of the transition band region in a manner that is strongly similar to that of Finite Impulse Response (FIR) filter design [20]. In fact, the development of MIMO beampattern synthesis methods has largely followed a parrallel path to that of optimal FIR filter design. Work by [2, 8] developed a series of algorithms that produced minimum mean-square error (MMSE) and minimax fits to desired beampatterns which are analogous to the efforts of [21, 22] that developed numerical optimization techniques for FIR filters. Work by [23] synthesized MIMO beampatterns using a Fourier sampling technique that is largely reminiscent of the work by Rabiner et. al. [24] for digital filter design. There are also several efforts that focus on beampattern synthesis with passband and stopband ripple control [25, 26, 27, 28] which mirrors some of the early attempts at equiripple FIR filter design [29]. Interestingly, [25] noted a relationship between the ripple levels, the length of the array, and the transition bandwidth of the beampattern that closely resembles the equiripple approximations developed by James Kaiser [30].
@@ II-A The MIMO Waveform Set
where $\varphi_{m}\left[n\right]$ is the $m^{\text{th}}$ waveform’s instantaneous phase and the $\sqrt{E/MN}$ term normalizes each waveform to have a total energy of $E/M$. This paper utilizes the MTSFM waveform model in [33] whose phase modulation function is a finite Fourier series
@@ II-B The MIMO Beampattern and Correlation Matrix
$\mathbf{X}=\left[x_{1}\left[n\right],x_{2}\left[n\right],\dots,x_{M}\left[n\right]\right]^{T}\in\mathcal{C}^{M\times N}.$ (3)
The narrowband MIMO transmit beampattern is a normalized power density across angle $\theta$ expressed as
$P\left(\theta\right)=\mathbf{a}^{H}\left(\theta\right)\mathbf{R}\mathbf{a}\left(\theta\right)\geq 0,~{}-\frac{\pi}{2}\leq\theta\leq\frac{\pi}{2}$ (4)
where $\mathbf{a}\left(\theta\right)$ is the transmit array steering vector
$\mathbf{a}\left(\theta\right)=\left[1,~{}e^{j\pi\sin\theta},~{}\dots,~{}e^{j\pi\left(M-1\right)\sin\theta}\right]^{\text{T}}.$ (5)
@@
$\tilde{\mathbf{X}}=\sum_{q=1}^{Q}w_{q}x_{q}=\mathbf{W}\mathbf{X}$ (7)
where $Q\leq M$ and $\mathbf{W}$ is the $M\times Q$ weights matrix whose $q^{\text{th}}$ column is $w_{q}$. The MIMO beampattern using (7) is now
$P\left(\theta\right)$ $=\mathbf{a}^{H}\left(\theta\right)\mathbf{R_{W}}\mathbf{a}\left(\theta\right)$
$=\mathbf{a}^{H}\left(\theta\right)\mathbf{W}\mathbf{X}\mathbf{X}^{H}\mathbf{W}^{H}\mathbf{a}\left(\theta\right)$ (8)
@@ III-A Problem Definition
This letter specifically focuses on MIMO beampatterns that are even-symmetric in angle for mathematical simplicity. However, we note that this method can be readily extended to arbitrary symmetry. As shown in Figure 1, the equiripple MIMO beampattern synthesis problem is essentially the same as the equiripple FIR filter design problem. There is a passband region $|u=\pi\sin\theta|\leq u_{p}$ across which the beampattern is flat at some power level $P_{0}$ with ripple $\delta_{p}=\delta$. There is also a well defined transition band $\Delta u=u_{s}-u_{p}$. What is different is that the stopband region $|u|\geq u_{s}$ has some non-zero level $\epsilon_{0}$ that must be greater than the stopband ripple $\delta_{s}=\delta_{p}=\delta$. This ensures that $P\left(u\right)$ is nonnegative for all $u$. Correspondingly, setting $\epsilon_{0}=\delta$ establishes the minimum achievable peak sidelobe height of the beampattern to be $2\delta$.
@@
$P\left(u\right)=\mathbf{a}\left(u\right)^{H}\mathbf{R}\mathbf{a}\left(u\right)=\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\mathbf{R}_{m,m^{\prime}}e^{-j\left(m-m^{\prime}\right)u}$ (10)
@@
The expression for the MIMO beampattern in (12) essentially possesses the form of a Type-I FIR filter frequency response [30]. Finding the coefficients $r_{\ell}$ to synthesize an equiripple beampattern can then be achieved using the Parks-McClellan algorithm [31]. Nonnegative trigonometric polynomials have been used in the design of FIR filters [37, 38] and appears to have only been briefly explored for MIMO beampattern synthesis [39, 40]. To the best of the authors’ knowledge, applying Chebyshev approximation to (12) to synthesize equiripple MIMO beampatterns appears to be novel.
The Chebyshev coefficients in (12) not only produces the desired equiripple MIMO beampattern, but also provides insight into the structure of the waveform correlation matrix that realizes it. Perhaps the simplest matrix construction is a Toeplitz structured matrix
@@ III-B Some Illustrative Design Examples
This section provides two illustrative design examples to demonstrate the equiripple MIMO beampattern synthesis method using Chebyshev approximation. For both design examples, MTSFM waveforms with a time-bandwidth product of $T\Delta f=64$ are utilized as in [34] where $T$ is the waveform duration in seconds and $\Delta f$ is the waveform’s swept bandwidth in Hz. The waveform time-series and are sampled at a rate $f_{s}=5\Delta f$ and therefore possess $N=Tf_{s}+1=5T\Delta f+1=321$ samples each.
@@ III-B1 ULA with $M=10$ Elements with Unit Elemental Power
The first design example utilizes a ULA with $M=10$ elements each transmitting equal power. The desired beampattern has unity gain (i.e $P_{0}=1$) in the passband defined as $|u|\leq u_{s}=0.2\pi$. The stopband is defined as $u_{s}=0.4\pi\leq|u|\leq\pi$ with a constant power level of $\epsilon_{0}=0.05$. Kaiser’s formula [30] dictates that the minimum ripple achievable is $\delta\approx 0.0112$. Thus, the peak sidelobe level in the stopband will be $\epsilon_{0}+\delta\approx 0.0612$ or $-12.1$ dB. The Fourier cosine coefficients $r_{\ell}$ from (12) for this equiripple beampattern design problem are found using the Parks-McClellelan algorithm. The corresponding correlation matrix can in fact be represented in the Toeplitz form according to (13) since it is positive semi-definite.
In order to find a MTSFM waveform set that realizes this beampattern, we do not require that its correlation matrix exactly fits the Toeplitz structure. Rather, only the Fourier cosine coefficients $r_{\ell}$ must match. This is achieved by solving the following constrained waveform optimization problem
@@
where $r_{\ell}$ are the Fourier cosine coefficients found using the Parks-McClellan algorithm, $\hat{r}_{\ell}\{\alpha_{m,p}\}$ are the coefficients resulting from the respective diagonals of the MTSFM waveform set’s correlation matrix $\mathbf{R}\{\alpha_{m,p}\}$, $\beta_{rms}^{2}\left(\{\alpha_{m,p}^{\left(i\right)}\}\right)$ is the Root Mean Square (RMS) bandwidth [42] of each MTSFM waveform in the set at the $i^{\text{th}}$ iteration of the routine, and $10<\mu<$ 1 is a unitless constant. The RMS bandwidth constraint, as discussed in [33], ensures that each MTSFM waveform’s spectral extent stays within some tolerance $\mu$ of the initial waveforms passed to (14) thus preserving its compact spectral shape [34]. The optimization routine described in (14) is solved using MATLAB’s optimization toolbox [43].
@@
The resulting equiripple beampatterns from the Toeplitz structured correlation matrix and the MTSFM waveform set are shown in panel (a) of Figure 2. Also shown in the figure are the Toeplitz and MTSFM correlation matrices that realize this equiripple beampattern. As can be seen in the figure, both equiripple beampatterns are essentially identical with a passband/stopband ripple of $\delta\approx 0.0112$ resulting in a peak sidelobe of roughly $-12.1$ dB. Note that while both beampatterns are identical, their respective correlation matrices possess different values for their off-diagonal entries. However, their summations along the $\ell^{\text{th}}$ order diagonals are identical, which is a consequence of (12).
@@ III-B2 ULA with $M=20$ Elements using TBP
The second design example utilizes a ULA with $M=20$ elements. Again, the desired beampattern has a unity gain $P_{0}=1$ in the passband defined as $|u|\leq u_{p}=0.2\pi$. The stopband is defined as $u_{s}=0.4\pi\leq|u|\leq\pi$ with a constant power level of $\epsilon_{0}=0.000339$ which is equal to the minimum ripple achievable per Kaiser’s formula [30]. Thus, the peak sidelobe level in the stopband is $\epsilon_{0}+\delta=0.000678$ or $-31.7$ dB. Once again, the coefficients $r_{\ell}$ from (12) are found using the Parks-McClellelan algorithm. However, the resulting Toeplitz structured correlation matrix is not positive semi-definite. Therefore, this example will use TBP with nonuniformly weighted elements that transmit orthogonal MTSFM waveforms to realize the equiripple MIMO beampattern. Finding a positive semi-definite correlation matrix that satisfies (12) is efficiently solved using the algorithm specified in (2.14) of [32] which also provides an efficient implementation using the CVX software package [44]. As was noted in [36], even small but nonzero correlation values between any two waveforms in the set can introduce perturbations in the TBP process that degrades the ideal MIMO beampattern shape. To minimize this effect, the MTSFM waveform set is optimized to be as nearly orthogonal as is practical with finite time-bandwidth product waveforms. This is achieved by solving the following constrained optimization problem.

### v2

@@ I Introduction
The majority of the efforts in the literature focus on designing beampatterns with a nominally flat passband region, a narrow transition band region, and a stopband region where almost no power is transmitted. As a consequence of the Gibbs phenonmenon, the beampattern’s passband & stopband regions possess ripples. The magnitude of these ripples appears to be controlled by the width of the transition band region in a manner that is strongly similar to that of Finite Impulse Response (FIR) filter design [20]. In fact, the development of MIMO beampattern synthesis methods has largely followed a parrallel path to that of optimal FIR filter design. Work by [2, 8] developed a series of algorithms that produced minimum mean-square error (MMSE) and minimax fits to desired beampatterns which are analogous to the efforts of [21, 22] that developed numerical optimization techniques for FIR filters. Work by [23] synthesized MIMO beampatterns using a Fourier sampling technique that is largely reminiscent of the work by Rabiner et. al. [24] for digital filter design. There are also several efforts that focus on beampattern synthesis with passband and stopband ripple control [25, 26, 27, 28] which mirrors some of the early attempts at equiripple FIR filter design [29]. Interestingly, [25] noted a relationship between the ripple levels, the length of the array, and the transition bandwidth of the beampattern that closely resembles the equiripple approximations developed by James Kaiser [30].
@@ II-A The MIMO Waveform Set
where $\varphi_{m}\left[n\right]$ is the $m^{\text{th}}$ waveform’s instantaneous phase and the $\sqrt{E/MN}$ term normalizes each waveform to have a total energy of $E/M$. We express (1) in vector form as $\mathbf{x}_{m}\in\mathcal{C}^{1\times N}$. This paper utilizes the MTSFM waveform model in [33] whose phase modulation function is a finite Fourier sine series
@@ II-B The MIMO Beampattern and Correlation Matrix
$\mathbf{X}=\left[\mathbf{x}_{1}^{T},\mathbf{x}_{2}^{T},\dots,\mathbf{x}_{M}^{T}\right]^{T}\in\mathcal{C}^{M\times N}.$ (3)
The narrowband MIMO transmit beampattern is a normalized power density across the spatial angle $u=\pi\sin\theta$
$P\left(u\right)=\mathbf{a}^{H}\left(u\right)\mathbf{R}\mathbf{a}\left(u\right)\geq 0,~{}-\pi\leq u\leq\pi$ (4)
where $\mathbf{a}\left(u\right)$ is the $M\times 1$ transmit array steering vector
$\mathbf{a}\left(u\right)=\left[1,~{}e^{ju},~{}\dots,~{}e^{j\left(M-1\right)u}\right]^{\text{T}}.$ (5)
@@
$\tilde{\mathbf{X}}=\sum_{m=1}^{M}\mathbf{w}_{m}\mathbf{x}_{m}=\mathbf{W}\mathbf{X}$ (7)
where $\mathbf{W}$ is the $M\times M$ weights matrix whose $m^{\text{th}}$ column is $\mathbf{w}_{m}$. The MIMO beampattern using (7) is now
$P\left(u\right)=\mathbf{a}^{H}\left(u\right)\mathbf{R_{W}}\mathbf{a}\left(u\right)=\mathbf{a}^{H}\left(u\right)\mathbf{W}\mathbf{W}^{H}\mathbf{a}\left(u\right)$ (8)
@@ III-A Problem Definition
This letter specifically focuses on MIMO beampatterns that are even-symmetric in angle for mathematical simplicity. However, we note that this method can be readily extended to arbitrary symmetry. As shown in Figure 1, the equiripple MIMO beampattern synthesis problem is essentially the same as the equiripple FIR filter design problem. There is a passband region $|u|\leq u_{p}$ across which the beampattern is flat at some power level $P_{0}$ with ripple $\delta_{p}=\delta$. There is also a well defined transition band $\Delta u=u_{s}-u_{p}$. What is different is that the stopband region $|u|\geq u_{s}$ has some non-zero level $\epsilon_{0}$ that must be greater than the stopband ripple $\delta_{s}=\delta_{p}=\delta$. This ensures that $P\left(u\right)$ is nonnegative for all $u$. Correspondingly, setting $\epsilon_{0}=\delta$ establishes the minimum achievable peak sidelobe height of the beampattern to be $2\delta$.
@@
$P\left(u\right)=\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\mathbf{R}_{m,m^{\prime}}e^{-j\left(m-m^{\prime}\right)u}$ (10)
@@
The expression for the MIMO beampattern in (12) essentially possesses the form of a Type-I FIR filter frequency response [30]. Finding the coefficients $r_{\ell}$ to synthesize an equiripple beampattern can then be achieved using the Parks-McClellan algorithm [31]. Choosing the necessary number of sensors $M$ to realize an equiripple beampattern shape with a desired $\Delta u$ and $\delta$ directly follows from a modified version of Kaiser’s formula for FIR filters [30]
$M\approx\dfrac{-20\log_{10}\left(\delta\right)-13}{14.6\Delta u}+1.$ (13)
Nonnegative trigonometric polynomials have been used in the design of FIR filters [37, 38] and appears to have only been briefly explored for MIMO beampattern synthesis [39, 40]. To the best of the authors’ knowledge, applying Chebyshev approximation to (12) to synthesize equiripple MIMO beampatterns appears to be novel.
The Chebyshev coefficients in (12) not only produce the desired equiripple MIMO beampattern, but it also provides insight into the structure of the waveform correlation matrix that realizes it. Perhaps the simplest matrix construction is a Toeplitz structured matrix
@@ III-B Some Illustrative Design Examples
This section provides two illustrative design examples to demonstrate the equiripple MIMO beampattern synthesis method using Chebyshev approximation. For both design examples, MTSFM waveforms with a time-bandwidth product of $T\Delta f=64$ are utilized as in [34] where $T$ is the waveform duration and $\Delta f$ is the waveform’s swept bandwidth. The number of harmonics $P$ in each MTSFM waveform’s instantaneous phase (2) were chosen to be $P=\left\lceil T\Delta f/2\right\rceil=32$ to ensure a compact spectral shape [33, 34]. The waveform time-series are sampled at a rate $f_{s}=5\Delta f$ and therefore possess $N=Tf_{s}+1=5T\Delta f+1=321$ samples each.
@@ III-B1 ULA with $M=10$ Elements with Unit Elemental Power
The first design example utilizes a ULA with $M=10$ elements each transmitting equal power. The desired beampattern has unity gain (i.e $P_{0}=1$) in the passband defined as $|u|\leq u_{p}=0.2\pi$. The stopband is defined as $u_{s}=0.4\pi\leq|u|\leq\pi$ with a constant power level of $\epsilon_{0}=0.05$. Equation (13) dictates that the minimum ripple achievable is $\delta\approx 0.0112$. The peak sidelobe level in the stopband is thus $\epsilon_{0}+\delta\approx 0.0612$ or $-12.1$ dB. The Fourier cosine coefficients $r_{\ell}$ from (12) for this equiripple beampattern design problem are found using the Parks-McClellelan algorithm. The corresponding correlation matrix resulting from (12) can be represented in the Toeplitz form according to (14) since it is positive semi-definite.
The Fourier cosine coefficients of the MTSFM waveform set that realizes this equiripple beampattern must match those found using the Parks-McClellan agorithm. This is achieved by solving the constrained waveform optimization problem
@@
where $r_{\ell}$ are the Fourier cosine coefficients found using the Parks-McClellan algorithm, $\hat{r}_{\ell}\{\alpha_{m,p}\}$ are the coefficients resulting from the respective diagonals of the MTSFM waveform set’s correlation matrix $\mathbf{R}\{\alpha_{m,p}\}$, $\beta_{rms}^{2}\left(\{\alpha_{m,p}^{\left(i\right)}\}\right)$ is the Root Mean Square (RMS) bandwidth [42] of each MTSFM waveform in the set at the $i^{\text{th}}$ iteration of the routine, and $10<\mu<$ 1 is a unitless constant. The RMS bandwidth constraint, as discussed in [33], ensures that each MTSFM waveform’s spectral extent stays within some tolerance $\mu$ of the initial waveforms passed to (15) thus preserving its compact spectral shape [34]. The optimization routine (15) is solved using MATLAB’s optimization toolbox [43].
@@
The resulting equiripple beampatterns from the Toeplitz structured correlation matrix and the MTSFM waveform set are shown in panel (a) of Figure 2. Also shown in the figure are the Toeplitz and MTSFM correlation matrices that realize this equiripple beampattern. As can be seen in the figure, both equiripple beampatterns are identical to machine precision with a passband/stopband ripple of $\delta\approx 0.0112$ resulting in a peak sidelobe of roughly $-12.1$ dB. Note that while both beampatterns are identical, their respective correlation matrices possess different values for their off-diagonal entries. However, their summations along the $\ell^{\text{th}}$ order diagonals are identical, which is a consequence of (12).
@@ III-B2 ULA with $M=20$ Elements using TBP
The second design example utilizes a ULA with $M=20$ elements. Again, the desired beampattern has a unity gain $P_{0}=1$ in the passband defined as $|u|\leq u_{p}=0.2\pi$. The stopband is defined as $u_{s}=0.4\pi\leq|u|\leq\pi$ with a constant power level of $\epsilon_{0}=\delta=0.000339$ derived via (13). The resulting peak sidelobe level in the stopband is thus $\epsilon_{0}+\delta=0.000678$ or $-31.7$ dB. Once again, the coefficients $r_{\ell}$ from (12) are found using the Parks-McClellelan algorithm. However, the resulting Toeplitz structured correlation matrix is not positive semi-definite. Heuristically, the authors found that for beampattern designs with equal element power and sidelobe levels below $\approx-13.2$ dB, the Toeplitz structured matrix was no longer positive semi-definite. Therefore, this example will use TBP with nonuniformly weighted elements that transmit orthogonal MTSFM waveforms to realize the equiripple MIMO beampattern. Finding a positive semi-definite correlation matrix that satisfies (12) is efficiently solved using the algorithm specified in (2.14) of [32] which also provides an efficient implementation using the CVX software package [44]. As was noted in [36], even small but nonzero correlation values between any two waveforms in the set can introduce perturbations in the TBP process that degrades the ideal MIMO beampattern shape. To minimize this effect, the MTSFM waveform set is optimized to be as nearly orthogonal as is practical with finite time-bandwidth product waveforms. This is achieved by solving the following constrained optimization problem